Erin is squaring numbers. She says that it is possible to get an answer that is smaller than the number she started with. Show how erin is correct.
Answer by nerdybill(7384)

If you square a number that is less than 1.
For example, say your number was 1/2
the square of 1/2 is (1/2)^2 = (1/2)(1/2) = 1/4